OpenStudy (anonymous):

Simplify by taking roots of the numerator and the denominator. Assume that all expressions under radicals represent positive numbers. 3√125x^11/y^3

OpenStudy (anonymous):

I got 457763671875√5x11/2 / y^3

OpenStudy (anonymous):

the cube root of y^3 is y the cubed root of 125 is 5 so you have \[\frac{5\sqrt{x^{11}}}{y}\] as a start

OpenStudy (anonymous):

then for \[\sqrt[3]{x^{11}}\] you think "3 goes in to 11 3 times with a remainder of 2, so x^3 comes out and x^2 stays in for a "final answer" of \[\frac{5x^3\sqrt[3]{x^2}}{y}\]

OpenStudy (anonymous):

is that a 3 ??

jimthompson5910 (jim_thompson5910):

Here is what satellite wrote in his last post but it's magnified a bit \[\LARGE \frac{5x^3\sqrt[3]{x^2}}{y}\]

OpenStudy (anonymous):

Thank you